OpenStudy (anonymous):

Crane A can unload a ship in 10 hours and crane B can unload it in 14 h. How long will it take the two cranes to unload the ship working together? What is the correct equation to solve this problem

OpenStudy (anonymous):

Hahah hella funny im doing this in math 120 hold on

OpenStudy (anonymous):

mine bellow is the same i just cant solve

OpenStudy (anonymous):

a*10=1 b*14=1 a = 1/10 b = 1/14

OpenStudy (anonymous):

( 1/10 + 1/14 ) X = 1 (job)

OpenStudy (anonymous):

LCD is 140, so its 14 + 10 = 140

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

24 X = 140 x = 140/24!

OpenStudy (anonymous):

almost proud of myself hahah, i got one like that but its using water and stuff. Cant solve : /

OpenStudy (anonymous):

Probally wanna simplify that 140/24

OpenStudy (anonymous):

Okay. Here is another one\[\frac{ a }{ 4 }-9=5\]

OpenStudy (anonymous):

just make it 1/4 - 9/1 = 5/1, find an LCD which is 4

OpenStudy (anonymous):

A. -16 B. -56 C. 56 D. 16